# Multi-Modal and Multi-Agent Systems Meet Rationality: A Survey

### v2

@@ Abstract
Rationality is the quality of being guided by reason, characterized by logical thinking and decision-making that align with evidence and logical rules. This quality is essential for effective problem-solving, as it ensures that solutions are well-founded and systematically derived. Despite the advancements of large language models (LLMs) in generating human-like text with remarkable accuracy, they present biases inherited from the training data, inconsistency across different contexts, and difficulty understanding complex scenarios involving multiple layers of context. Therefore, recent research attempts to leverage the strength of multiple agents working collaboratively with various types of data and tools for enhanced consistency and reliability. To that end, this paper aims to understand whether multi-modal and multi-agent systems are advancing toward rationality by surveying the state-of-the-art works, identifying advancements over single-agent and single-modal systems in terms of rationality, and discussing open problems and future directions. We maintain an open repository at https://github.com/bowen-upenn/MMMA_Rationality.
@@
Bowen Jiang1, Yangxinyu Xie1, 2, Xiaomeng Wang1, Weijie J. Su1 Camillo J. Taylor1, Tanwi Mallick2    University of Pennsylvania1    Argonne National Laboratory2 Philadelphia, PA, 19104, USA   Lemont, IL, 60439, USA {bwjiang@seas, xinyux@wharton, cjtaylor@seas}.upenn.edu, tmallick@anl.gov
@@ 1 Introduction
Large language models (LLMs) have demonstrated promising results across a broad spectrum of tasks, particularly in exhibiting capabilities that plausibly mimic human-like reasoning Wei et al. (2022); Yao et al. (2024); Besta et al. (2024); Shinn et al. (2024); Bubeck et al. (2023); Valmeekam et al. (2023); Prasad et al. (2023). These models leverage the richness of human language to abstract concepts, elaborate thinking process, comprehend complex user queries, and develop plans and solutions in decision-making scenarios. Despite these advances, recent research has revealed that even state-of-the-art LLMs exhibit various forms of irrational behaviors, such as the framing effect, certainty effect, overweighting bias, and conjunction fallacy Binz and Schulz (2023); Echterhoff et al. (2024); Mukherjee and Chang (2024); Macmillan-Scott and Musolesi (2024); Wang et al. (2024a); Suri et al. (2024). These biases significantly challenge the utility of LLMs in natural language processing research. For example, LLM-based evaluators, a popular choice for automated assessments for text generation, display cognitive biases against certain responses irrespective of their actual quality or relevance Stureborg et al. (2024); Koo et al. (2023). Irrationality and hallucinations  Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023) also undermine the practical deployment of LLMs in critical sectors like healthcare, finance, and legal services He et al. (2023); Li et al. (2023i); Kang and Liu (2023); Cheong et al. (2024), where reliability and consistency are paramount. The emerging concern about the factual accuracy and trustworthiness of LLMs highlighting an urgent need to develop better agents or agent systems Nakajima (2023); Gravitas (2023) with rational reasoning processes.
One possible reason for the LLMs’ irrational behaviors, as suggested by Bubeck et al. (2023) and Sun (2024), is the autoregressive nature of existing language models. This architecture doesn’t allow for an “internal scratchpad" beyond these models’ inner parametric representations of knowledge, causing them to fail to reason rationally when faced with problems that require more complex and iterative procedures. Thus, an important question emerges: How can we design an LLM-based agent capable of rational decision-making that can overcome these biases and inconsistencies?
@@
Recent advancements in multi-modal and multi-agent frameworks offer a promising direction to address this challenge, which leverage the expertise of different agents acting together towards a collective goal. Multi-modal foundation models Awadalla et al. (2023); Liu et al. (2023b); Wang et al. (2023d); OpenAI (2023); Reid et al. (2024) enhance reasoning by grounding decisions in a broader sensory context, akin to how human brains integrate rich sensory inputs to form a more holistic base of knowledge. Meanwhile, multi-agent systems introduce mechanisms such as consensus, debate, and self-consistency Du et al. (2023); Liang et al. (2023); Talebirad and Nadiri (2023); Madaan et al. (2024); Cohen et al. (2023); Shinn et al. (2024); Mohtashami et al. (2023), which allow for more refined and reliable output through collaborative interaction among multiple instances. Each agent is specialized in different domains and offers its unique perspectives, simulating the dynamics of discussion in human societies. Multi-agent systems can also incorporate multi-modal agents and agents specialized in querying external knowledge sources or tools Lewis et al. (2020); Schick et al. (2024); Tang et al. (2023); Pan et al. (2024) to overcome hallucinations, ensuring that their results are more robust, deterministic, and trustworthy, thus significantly improving the quality of the generated responses towards rationality.
@@
This survey provides a unique lens to interpret the underlying motivations behind current multi-modal and/or multi-agent systems. Drawing from cognitive science, we first delineate four fundamental requirements for rational thinking in Section 2, and then Section 4 discusses how research fields within the multi-modality and multi-agents literature are progressing towards rationality by inherently improving these criteria. We posit that such advancements are bridging the gap between the performance of these systems and the expectations for a rational thinker, in contrast to traditional single-agent language-only models. Furthermore, Section 5 highlights the lack of sufficient evaluation metrics and benchmarks in the existing literature to adequately measure the rationality of LLMs or agent systems. We hope this survey can inspire further research at the intersection between agent systems and cognitive science.
@@ 2 Defining Rationality
Figure 1: The evolutionary tree of multi-agent and/or multi-modal systems related to the four axioms of rationality. Many proposed approaches strive to address multiple axioms simultaneously. Bold fonts are used to mark works that involve multi-modalities. This tree also includes foundational works to provide a clearer reference of time.
A rational agent should avoid reaching contradictory conclusions in decision making processes, respecting the physical and factual reality of the world in which it operates. Thus, drawing on foundational works in rational decision-making (Tversky and Kahneman, 1988; Hastie and Dawes, 2009; Eisenführ et al., 2010), this section adopts an axiomatic approach to define rationality, presenting four substantive axioms that we expect a rational agent or agent systems to fulfill:
@@
The decision of a rational agent is grounded on the physical and factual reality. For example, a flight booking agent must accurately retrieve available airports without fabricating nonexistent ones, and a video generation agent should adhere to the laws of physics in a world model.
Naturally, in order to make a sound decision, the agent must be able to integrate sufficient and accurate information from different sources and modalities grounded in reality without hallucination. While this requirement is generally not explicitly stated in the cognitive science literature when defining rationality, it is implicitly implied, as most humans have access to physical reality through multiple sensory signals.
@@
When comparing alternatives in a decision scenario, a rational agent can rank the options based on the current state and ultimately select the most preferred one based on the expected outcomes. This orderability consists of several key principles, including comparability, transitivity closure, solvability, etc. with detailed defined in Appexdix A. The orderability of preferences ensures the agent can make consistent and logical choices when faced with multiple alternatives. LLM-based evaluations heavily rely on this property, as discussed in Appendix B.
@@
The agent’s preference should not be influenced by information irrelevant to the decision problem at hand. LLMs have been shown to exhibit irrational behavior when presented with irrelevant context Shi et al. (2023); Wu et al. (2024); Liu et al. (2024d); Yoran et al. (2023), leading to confusion and suboptimal decisions. To ensure rationality, an agent must be able to identify and disregard irrelevant information, focusing solely on the factors that directly impact the decision-making processes.
@@ 3 Scope
Unlike existing surveys Han et al. (2024); Guo et al. (2024); Xie et al. (2024a); Durante et al. (2024); Cui et al. (2024); Xu et al. (2024b); Zhang et al. (2024a); Cheng et al. (2024); Li et al. (2024a) that focus on the components, structures, agent profiling, planning, communications, memories, and applications of multi-modal and/or multi-agent systems, this survey is the first to specifically examine the increasingly important relations between rationality and these multi-modal and multi-agent systems, exploring how they contribute to enhancing the rationality in decision-making processes.
@@
Rationality is also different from Theory of Mind (ToM) Apperly and Butterfill (2009); Nye et al. (2021); Oguntola et al. (2021); Hagendorff (2023); Li et al. (2023c); Sclar et al. (2023b); Kosinski (2023) in machine psychology. ToM refers to the model’s ability to understand that others’ mental states, beliefs, desires, emotions, and intentions may be different from its own.
@@ 4 Towards Rationality through Multi-Modal and Multi-Agent Systems
This section surveys recent advancements in multi-modal and multi-agent systems, categorized by their fields as depicted in Figure 1.
Each category of research in Figure 1, such as knowledge retrieval or neuro-symbolic reasoning, addresses one or more fundamental requirements for rational thinking. These rationality requirements are typically intertwined; therefore, an approach that enhances one aspect of rationality often inherently improves others simultaneously. Meanwhile, the overall goal of current multi-agent system in achieving rationality can usually be distilled into two key concepts: deliberation and abstraction. Deliberation encourages slower reasoning process such as brainstorming and reflection, while abstraction refers to boiling down the problem its into logical essence like calling APIs of tools or incorporating neuro-symbolic reasoning agents.
@@
Besides, dual-process theories Sun (2001); Evans (2003); Kahneman (2011); Sun (2024); Nye et al. (2021) in human cognition suggest that a single LLM, in analogy, might engage primarily in “System 1" thinking that is fast and automatic, but prone to biases. In contrast, “System 2" thinking is slow and rule-based (Kahneman, 2011), which is more reliable but requires more cognitive effort. This section will also discuss how multi-modal and multi-agent systems align with these theories.
4.1 Towards Grounding & Invariance through Multi-Modal Models
Multi-modal approaches aim to improve the information grounding across various channels, such as language and vision. By incorporating multi-modal agents, multi-agent systems can greatly expand their capabilities, enabling a richer, more accurate, and contextually aware interpretation of environment.
@@ 4.1 Towards Grounding & Invariance through Multi-Modal Models
Grounding an agent solely based on textual language can be challenging, as information can be represented much more efficiently through other sensory modes. For example, as a picture is worth a thousand words, recent advances in large vision-language pretraining have enabled LLMs with robust language comprehension capabilities to finally perceive the visual world. Multi-modal foundation models, including but not limited to CLIP Radford et al. (2021), VLBERT and ViLBERT Su et al. (2019); Lu et al. (2019), BLIP-2 Li et al. (2023e), (Open) Flamingo Alayrac et al. (2022); Awadalla et al. (2023), LLaVA Liu et al. (2024c, 2023b), CogVLM Wang et al. (2023d), MiniGPT-4 Zhu et al. (2023a), GPT-4 Vision OpenAI (2023) and GPT-4o OpenAI (2024), and Gemini 1.5 Pro Reid et al. (2024) serve as the cornerstones for multi-modal agent systems to ground knowledge in vision and beyond.
Alignment with Dual-Process Theories
Generating output texts from input images in these models requires only a single inference pass, which is quick and straightforward. This process bypasses the need for iterative reasoning or reflective steps, aligning single-model inference closely with the System 1 process of fast and automatic thinking.
The adaptation of Reinforcement Learning from Human Feedback (RLHF) Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022), a technique popularized in language-only models, in LLaVA-RLHF Sun et al. (2023b) demonstrates promising advancements in reducing hallucination from cross-modal misalignment. Furthermore, visual instruction-tuning Liu et al. (2024c); Dai et al. (2024); Bai et al. (2023); Wang et al. (2023c) enables advanced foundation models like LLaVA Liu et al. (2024c, 2023b), GPT-4 Vision OpenAI (2023), and Gemini 1.5 Pro Reid et al. (2024) to engage in more detailed instruction following, multi-round human-agent interactions, and collaborations with other agents, performing deliberate multimodal reasoning, tool using Schick et al. (2024), and strategic thinking to understand more engineered or context-dependent queries. This opens the possibility of subsequent research on the System 2 process in multi-modal foundation models.
@@
Multi-modalities help enhance the functionality of agent systems through more diverse information grounding. For example, Chain-of-Action Pan et al. (2024) advances the single-modal Search-in-the-Chain Xu et al. (2023) by supporting multi-modal data retrieval for faithful question answering. DoraemonGPT Yang et al. (2024) decomposes complex tasks into simpler ones toward understanding dynamic scenes, where multi-modal understanding is necessary for spatial-temporal videos analysis. RA-CM3 Yasunaga et al. (2022) augments baseline retrieval-augmented LLMs with raw multi-modal documents that include both images and texts, assuming that these two modalities can contextualize each other and make the documents more informative, leading to better generator performance. The multi-modal capabilities also allow HuggingGPT Shen et al. (2024b), Agent LUMOS Yin et al. (2023), ToolAlpaca Tang et al. (2023), and AssistGPT Gao et al. (2023b) to expand the scope of tasks they can address, including cooperation among specialized agents or tools capable of handling different information modalities.
Web agents are another example of how multi-modal agents surpass language-only ones. In agents like Pix2Act Shaw et al. (2024), WebGUM Furuta et al. (2023), CogAgentHong et al. (2023b), and SeeAct Zheng et al. (2024a), web navigation is grounded on graphical user interface (GUI) rather than solely on HTML texts Shen et al. (2024a); Yao et al. (2022a); Deng et al. (2024); Gur et al. (2023). This method of visual grounding offers higher information density compared to HTML codes that are usually lengthy, noisy, and sometimes even incomplete Zheng et al. (2024a). Supporting the importance of vision, ablation studies in WebGUM Furuta et al. (2023) also reports $5.5\%$ success rate improvement on the MiniWoB++ dataset Shi et al. (2017); Liu et al. (2018) by simply adding the image modality.
Large world models is an emerging and promising direction to reduce multimodal hallucinations. The notion is also mentioned in “Objective-driven AI" LeCun (2024), where agents have behavior driven by fulfilling objectives, i.e., drives, and they understand how the world works with common sense knowledge, beyond an auto-regressive generation. LeCun (2024) proposes the urgency for agents to learn to reason beyond feed-forward, i.e., the System 1 subconscious computation, and start making System 2 reasoning and planning on complicated actions to satisfy objectives with a grounding on world models. For example, Ghost-in-the-Minecraft Zhu et al. (2023b) and Voyager Wang et al. (2023a) have agents living in a well-defined game-world environment. JEPA LeCun (2022) creates a recurrent world model in an abstract representation space. Large World Model (LWM) Liu et al. (2024b) and Sora Brooks et al. (2024) develop insights from both textual knowledge and the world through video sequences. They both advance toward general-purpose simulators of the world, but still lack reliable physical engines for guaranteed grounding in real-world dynamics.
Invariance Across Modalities
Achieving representation invariance across modalities is another critical facet of rationality. Given adequate information grounding, agents should make consistent decisions across different modalities that share equivalent underlying logic.
Multi-modal foundation models are particularly adept at promoting invariance by processing multi-modal data in an unified representation. Specifically, their large-scale cross-modal pretraining stage seamlessly tokenizes both vision and language inputs into a joint hidden embedding space, learning cross-modal correlations through a data-driven approach. In other words, image tokens are simply regarded as a foreign language Wang et al. (2022a). Moreover, the cross-modal validation inherent in multi-modal foundation models allows for reconciliation of data from different modalities, closing their distance in the hidden embedding space Radford et al. (2021).
The concept of invariance is the cornerstone of Visual Question Answering (VQA) agents Chen et al. (2022); Jiang et al. (2024); Wang et al. (2023e); Yi et al. (2018); Wang et al. (2022a); Bao et al. (2022); Zhao and Xu (2023). On one hand, these agents must grasp the invariant semantics of any open-ended questions posed about images, maintaining consistency despite variations in wording, syntax, or language. On the other hand, within a multi-agent VQA system, visual agents can provide crucial verification and support for language-based reasoning Wang et al. (2023e); Jiang et al. (2024); Zhao and Xu (2023), while language queries can direct the attention of visual agents, based on a shared and invariant underlying knowledge across vision and language domains.
4.2 Towards Grounding through Knowledge Retrieval
@@
Bounded Rationality March and Simon (1958); Selten (1990) is a concept tailored to cognitively limited agents, suggesting that decision-making is limited by the resources available at hand, and any deviations from the optimal are primarily due to insufficient computational capacity and bounded working memory. In terms of LLMs, the parametric nature of their existing architecture Vaswani et al. (2017) fundamentally limits how much information they can hold. As a result, in the face of uncertainty, LLMs often hallucinate  Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023), generating outputs that are not supported by the factual reality of the environment. Retrieval-Augmented Generation (RAG) Lewis et al. (2020) marks a significant milestone in addressing such an inherent limitation of LLMs. Broadly speaking, RAG refers to any mechanism that provides external knowledge to the input context of an LLM and helps it deliver responses with up-to-date, factual, and grounded information.
A multi-modal and/or multi-agent system can include planning agents in its framework, which is akin to the System 2 process that can determine how and where to retrieve external knowledge, and what specific information to acquire. Additionally, the system can have summarizing agents that utilize retrieved knowledge to enrich the system’s language outputs with better factuality.
There are multiple works that construct large-scale knowledge graphs (KGs) Hogan et al. (2021) from real-world sources to effectively expand their working memory. Specifically, compared to language-only models, MAVEx Wu et al. (2022) improves system’s scores by $9.5\%$ compared to an image-only baseline through the integration of knowledge from ConceptNet Speer et al. (2017) and Wikipedia Wikipedia contributors (2004). It also improves the scores by $8.3\%$ by using the image modality for cross-modal validations with an oracle. Thanks to the external knowledge base, ReAct Yao et al. (2022b) reduces false positive rates from hallucination by $8.0\%$ compared to CoT Wei et al. (2022). CuriousLLM Yang and Zhu (2024) presents ablation studies showing the effectiveness of KGs on improving reasoning within the search process. MineDojo Fan et al. (2022) observes that internet-scale multi-modal knowledge allows models to significantly outperform all creative task baselines. Equipped with world knowledge, RA-CM3 Yasunaga et al. (2022) can finally generate faithful images from captions compared to CM3 Aghajanyan et al. (2022) and Stable Diffusion Rombach et al. (2022). CooperKGC Ye et al. (2023) enables multi-agent collaborations, leveraging knowledge bases of different experts. It finds that the incorporation of KGs improves F1 scores by $10.0$-$33.6\%$ across different backgrounds, and adding more collaboration rounds also enhance performance by about $10.0$-$30.0\%$. DoraemonGPT Yang et al. (2024) supports knowledge tools to assist the understanding of specialized video contents. SIRI Wang et al. (2023e) builds a multi-view knowledge base to increase the explainability of visual question answering. Grounding agents in external knowledge base also promotes more factual rationales and fewer hallucinations, especially in scientific and medical domains, exemplified by Chameleon Lu et al. (2024), Chain-of-Knowledge Li et al. (2023g), WildfireGPT Xie et al. (2024b), and Agent Hospital Li et al. (2024b). Chain-of-Knowledge Li et al. (2023g) even discovers that integrating multiple knowledge sources enhances performance by $2.1\%$ compared to using a single source in its experiments.
4.3 Towards Grounding & Invariance & Independence from Irrelevant Contexts through Tool Utilization
Information Grounding

### v3

@@ Abstract
Rationality is the quality of being guided by reason, characterized by decision-making aligned with evidence and logical rules. This quality is essential for effective problem-solving, as it ensures that solutions are well-founded and consistently derived. Despite the advancements of large language models (LLMs) in generating human-like texts with remarkable accuracy, they present limited knowledge space, inconsistency across contexts, and difficulty understanding complex scenarios. Therefore, recent research focuses on building multi-modal and multi-agent systems to achieve considerable progress with enhanced consistency and reliability, instead of relying on a single LLM as the sole planning or decision-making agent. To that end, this paper aims to understand whether multi-modal and multi-agent systems are advancing toward rationality by surveying the state-of-the-art works, identifying advancements over single-agent and single-modal systems in terms of rationality, and discussing open problems and future directions. We maintain an open repository at https://github.com/bowen-upenn/MMMA_Rationality.
@@
Bowen Jiang1, Yangxinyu Xie1, 2, Xiaomeng Wang1, Weijie J. Su1, Camillo J. Taylor1, Tanwi Mallick2    University of Pennsylvania1    Argonne National Laboratory2 Philadelphia, PA, 19104, USA   Lemont, IL, 60439, USA {bwjiang@seas, xinyux@wharton, xwang1@wharton}.upenn.edu {suw@wharton, cjtaylor@seas}.upenn.edu, tmallick@anl.gov
@@ 1 Introduction
Large language models (LLMs) have demonstrated promising results across a broad spectrum of tasks, particularly in exhibiting capabilities that plausibly mimic human-like reasoning Wei et al. (2022); Yao et al. (2024); Besta et al. (2024); Shinn et al. (2024); Bubeck et al. (2023); Valmeekam et al. (2023); Prasad et al. (2023). These models leverage the richness of human language to abstract concepts, elaborate thinking process, comprehend complex user queries, and develop plans and solutions in decision-making scenarios. Despite these advances, recent research has revealed that even state-of-the-art LLMs exhibit various forms of irrational behaviors, such as the framing effect, certainty effect, overweighting bias, and conjunction fallacy Binz and Schulz (2023); Echterhoff et al. (2024); Mukherjee and Chang (2024); Macmillan-Scott and Musolesi (2024); Wang et al. (2024a); Suri et al. (2024). Irrationality undermine the practical deployment of LLMs in critical sectors like healthcare, finance, and legal services He et al. (2023); Li et al. (2023h); Kang and Liu (2023); Cheong et al. (2024), where reliability and consistency are paramount. The emerging concern about the factual accuracy and trustworthiness of LLMs highlights an urgent need to develop better agents or agent systems Nakajima (2023); Gravitas (2023) with rational reasoning processes.
@@
A single LLM agent can fall into irrational behaviors because it cannot go beyond the language model’s inner parametric representations of textual knowledge, lacking the real-world grounding and feedback mechanisms necessary to develop rationality (Bubeck et al., 2023; Sun, 2024). In contrast, in real life scenarios, important decisions are rarely made by individuals on their own, and the complexity of problems often requires the collaboration of experts from different fields to ensure rationality (Eisenführ et al., 2010). In a similar vein, recent advancements in multi-modal and multi-agent frameworks leverage the expertise of different agents acting together towards a collective goal. Multi-modal foundation models Awadalla et al. (2023); Liu et al. (2023a); Wang et al. (2023c); OpenAI (2023); Reid et al. (2024) enhance reasoning by grounding decisions in a broader sensory context, akin to how human brains integrate rich sensory inputs to form a more holistic base of knowledge. Meanwhile, multi-agent systems introduce mechanisms such as consensus, debate, and self-consistency Du et al. (2023); Liang et al. (2023); Talebirad and Nadiri (2023); Madaan et al. (2024); Cohen et al. (2023); Shinn et al. (2024); Mohtashami et al. (2023) to allow for refined and reliable output through collaborative interactions. Such systems can also query external knowledge sources or tools Lewis et al. (2020); Schick et al. (2024); Tang et al. (2023); Pan et al. (2024) to augment their reasoning capabilities for rational decision making.
@@
This survey provides a unique lens to reinterpret the underlying motivations behind current multi-modal and/or multi-agent systems by drawing insights from cognitive science. In Section 2, we outline four essential requirements for rational decision making. Section 4 then examines how various research areas within the multi-modality and multi-agent literature are advancing towards rationality based on these criteria. We argue that these advancements surpass the limitations of single language-model agents and narrow the gap between the behavior of agent systems and the expectations for rational decision making. Lastly, Section 5 highlights the lack of sufficient evaluation metrics and benchmarks in the existing literature to adequately measure the rationality of LLMs or agent systems. We hope this survey can inspire further research at the intersection between agent systems and cognitive science.
@@ 2 Defining Rationality
Figure 1: The evolutionary tree of multi-agent and/or multi-modal systems related to the four axioms of rationality. Many proposed approaches strive to address multiple axioms simultaneously. The bold font marks works that involve multi-modalities. This tree also includes some foundational works to provide a clearer reference of time.
A rational agent should respect the reality of the world in which it operates and avoid reaching contradictory conclusions in decision-making processes. Drawing on foundational works in rational decision-making (Tversky and Kahneman, 1988; Hastie and Dawes, 2009; Eisenführ et al., 2010), this section adopts an axiomatic approach to define rationality, presenting four substantive axioms we expect a rational agent or agent systems to fulfill:
@@
The decision of a rational agent is grounded on the physical and factual reality. For example, a video generation agent should adhere to the laws of physics in a world model and a forecasting assistant ought to estimate likelihoods obeying the law of probability.
@@
When comparing alternatives in a decision scenario, a rational agent can rank the options based on the current state and ultimately select the most preferred one based on the expected outcomes. This orderability consists of several key principles, including comparability, transitivity closure, solvability, etc. with detailed defined in Appexdix A.
@@
The agent’s preference should not be influenced by information irrelevant to the decision-making problem at hand.
@@ 3 Scope
Unlike existing surveys Han et al. (2024); Guo et al. (2024); Xie et al. (2024a); Durante et al. (2024); Cui et al. (2024); Xu et al. (2024b); Zhang et al. (2024a); Cheng et al. (2024); Li et al. (2024a) that focus on the components, structures, agent profiling, planning, communications, memories, and applications of multi-modal and/or multi-agent systems, this survey is the first to specifically examine the increasingly important relations between rationality and these multi-modal and multi-agent systems, exploring how they contribute to enhancing the robustness in decision-making processes.
@@ 4 Towards Rationality through Multi-Modal and Multi-Agent Systems
This section surveys recent advancements in multi-modal and multi-agent systems under different research categories as depicted in Figure 1. Each category, such as knowledge retrieval or neuro-symbolic reasoning, addresses one or more fundamental requirements for rational thinking. These rationality requirements are typically intertwined: an approach that enhances one aspect of rationality often inherently improves others simultaneously. Meanwhile, the overall mechanism of current multi-agent system in achieving rationality can categorized into two key concepts: deliberation and abstraction. Deliberation encourages a slower, iterative reasoning process, while abstraction refers to abstracting the problem into its logical essence.
@@
4.1 Towards Grounding & Invariance
Multi-modal approaches aim to improve information grounding across various channels, such as language, vision, and beyond. By incorporating multi-modal agents, multi-agent systems can greatly expand their capabilities, enabling a richer, more accurate, and contextually aware interpretation of the environment.
@@ 4.1 Towards Grounding & Invariance
Grounding an agent solely based on textual language can be challenging, as information can be represented much more efficiently through other sensory modes. As a picture is worth a thousand words, recent advances in large vision-language pretraining have enabled LLMs with robust language comprehension capabilities to finally perceive the visual world. Multi-modal foundation models, including but not limited to CLIP Radford et al. (2021), VLBERT and ViLBERT Su et al. (2019); Lu et al. (2019), BLIP-2 Li et al. (2023d), (Open) Flamingo Alayrac et al. (2022); Awadalla et al. (2023), LLaVA Liu et al. (2024c, 2023a), CogVLM Wang et al. (2023c), MiniGPT-4 Zhu et al. (2023a), GPT-4 Vision OpenAI (2023) and GPT-4o OpenAI (2024), and Gemini 1.5 Pro Reid et al. (2024) serve as the cornerstones for multi-modal agent systems to ground knowledge in vision and beyond.
Invariance Across Modalities
Achieving representation invariance across modalities is critical: given adequate information grounding, agents should make consistent decisions across different modalities that share equivalent underlying logic. Multi-modal foundation models are particularly adept at promoting invariance by processing multi-modal data in an unified representation. Specifically, their large-scale cross-modal pretraining stage seamlessly tokenizes both vision and language inputs into a joint hidden embedding space, learning cross-modal correlations through a data-driven approach. In other words, image tokens are simply regarded as a foreign language Wang et al. (2022a). Moreover, the cross-modal validation inherent in multi-modal foundation models allows for reconciliation of data from different modalities, closing their distance in the hidden embedding space Radford et al. (2021).
The concept of invariance is the cornerstone of Visual Question Answering (VQA) agents Chen et al. (2022); Jiang et al. (2024b); Wang et al. (2023d); Yi et al. (2018); Wang et al. (2022a); Bao et al. (2022); Zhao and Xu (2023). On one hand, these agents must grasp the invariant semantics of any open-ended questions posed about images, maintaining consistency despite variations in wording, syntax, or language. On the other hand, within a multi-agent VQA system, visual agents can provide crucial verification and support for language-based reasoning Wang et al. (2023d); Jiang et al. (2024b); Zhao and Xu (2023), while language queries can direct the attention of visual agents, based on a shared and invariant underlying knowledge across vision and language domains.
@@
Multi-modalities help enhance the functionality of agent systems through more diverse information grounding. Web agents are a quintessential example of how multi-modal agents surpass language-only ones. Because HTML code is often lengthy, contains irrelevant information, and may be incomplete Zheng et al. (2024a), web navigation grounded on the graphical user interface (GUI) offers higher information density compared to solely HTML codes. As a result, using visual cues from the GUI leads to improved navigation performance Shen et al. (2024a); Yao et al. (2022a); Deng et al. (2024); Gur et al. (2023). Multi-modalities also help enhance the functionality of agent systems through more diverse information grounding. For example, RA-CM3 Yasunaga et al. (2022) augments baseline retrieval-augmented LLMs with raw multi-modal documents that include both images and texts, assuming that these two modalities can contextualize each other and make the documents more informative, leading to better generator performance. For other examples, we refer the reader to Appendix B.
Knowledge Retrieval & Tool Usage
@@
Bounded Rationality March and Simon (1958); Selten (1990) is a concept tailored to cognitively limited agents, suggesting that decision-making is limited by the resources available at hand, and any deviations from the optimal are primarily due to insufficient computational capacity and bounded working memory. In terms of LLMs, the parametric nature of their existing architecture Vaswani et al. (2017) fundamentally limits how much information they can hold. As a result, in the face of uncertainty, LLMs often hallucinate  Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023), generating outputs that are not supported by the factual reality of the environment. Retrieval-Augmented Generation (RAG) Lewis et al. (2020) marks a significant milestone in addressing such an inherent limitation of LLMs. Broadly speaking, RAG refers to any mechanism that provides external knowledge to the input context of an LLM and helps it deliver responses with up-to-date, factual, and grounded information , especially in scientific and medical domains. Examples include Chameleon Lu et al. (2024), Chain-of-Knowledge Li et al. (2023f), WildfireGPT Xie et al. (2024b), and Agent Hospital Li et al. (2024b). Specifically, Chain-of-Knowledge Li et al. (2023f) even discovers that integrating multiple knowledge sources enhances performance by $2.1\%$ compared to using a single source in its experiments.
Another line of systems construct large-scale knowledge graphs (KGs) Hogan et al. (2021) from real-world sources to effectively expand their working memory and improve their task performance. Specifically, compared to language-only models, MAVEx Wu et al. (2022) improves system’s scores by $9.5\%$ compared to an image-only baseline through the integration of knowledge from ConceptNet Speer et al. (2017) and Wikipedia Wikipedia contributors (2004). It also improves the scores by $8.3\%$ by using the image modality for cross-modal validations with an oracle. Thanks to the external knowledge base, ReAct Yao et al. (2022b) reduces false positive rates from hallucination by $8.0\%$ compared to CoT Wei et al. (2022). For more examples, see Appendix C.